OpenStudy (anonymous):

HELP !!!! Use the following set of data for parts : 101, 102, 100, 100, 109, 110, 109, 108, 109 d. Find the variance _________ e. Find the standard deviation __________

OpenStudy (anonymous):

101-105.3^2= 18.49 102-105.3^2= -10.89 100-105.3^2= 28.09 100-105.3^2= 28.09 109-105.3^2= 13.69 110-105.3^2=22.09 109-105.3^2= 13.69 108-105.3= 7.29 109-105.3^2= 13.69

OpenStudy (anonymous):

156.01